# DRAGON: Distributional Rewards Optimize Diffusion Generative Models

### v1

@@ 5.1 Models, Datasets, Training Settings, and Evaluation Metrics
Baseline model and pre-training. We use the base diffusion model from Presto (Novack et al., 2025) to generate 32-second single-channel (mono) 44.1kHz audio. It includes a latent-space score-prediction de-noising module (Rombach et al., 2022; Karras et al., 2022; Song et al., 2021b) based on DiT-XL (Peebles & Xie, 2023) that takes in the noise level and a text embedding as conditioning signals, a convolutional variational autoencoder (VAE) that converts audio to and from the diffusion latent space (Kumar et al., 2023), and a FLAN-T5-based text encoder (Chung et al., 2024a). The baseline model is pre-trained with diffusion loss to convergence on a 3600-hour instrumental music dataset with musical-metadata-grounded synthetic captions, which we call the Adobe Licensed Instrumental Music dataset (ALIM). Our inference uses 40 diffusion steps with the second-order DPM sampler (Lu et al., 2022) with CFG++ ($w=0.8$) enabled in selected time steps (Chung et al., 2024b). Please see Appendix E for details.
Training and evaluation prompts. We use ALIM training prompts (same setting as in pre-training) for DRAGON fine-tuning. Our evaluation uses a combination of the captions in an independent ALIM test split (800 pieces), the captions in a non-vocal Song Describer subset (Manco et al., 2023) (585 pieces, abbreviated as SDNV), and the real-world user prompts in the DMA dataset (800 pieces). Unless specified otherwise, all evaluation metrics are computed with generations from these 2,185 prompts (one generation per prompt).
@@
Correlation measures. In addition to evaluating and comparing DRAGON models, we offer correlation analyses between various reward signals. We quantify correlation with overall Pearson correlation (PLCC) and per-prompt Spearman’s rank correlation (SRCC) in Section B.2.
FAD encoders. The per-song and dataset FAD use ALIM and SDNV as reference statistics. DRAGON training uses ALIM’s training split statistics to compute FAD, while evaluation uses the test split. We consider CLAP and the diffusion VAE as FAD encoders, with the former known to provide high-quality semantic embeddings (Gui et al., 2024). To account for any VAE reconstruction inaccuracies, we decode the generated VAE embeddings to audio and back. Compared with CLAP, the VAE encoder produces considerably more embeddings per song (1,836 versus 9) but in a much lower dimension (32 versus 512).
CLAP embedding setting. We use the LAION-CLAP checkpoint that specializes in music (Wu* et al., 2023; Chen et al., 2022). Since CLAP takes in 10-second 48kHz audio whereas our generations are 32-second 44.1kHz, pre/post-processing is required. We focus on two settings: MA and FADTK. Our MA setting is optimized for our music aesthetics predictor with ablation studies in Section B.2 while our FADTK setting follows (Gui et al., 2024). More details and a comparison between the two settings are in Section C.2.
Open-source comparison. We compare DRAGON with open-source models such as Stable Audio (Evans et al., 2024a) and MusicGen (Copet et al., 2023) in Appendix D.
We first use DRAGON to optimize instance-wise (reference-free) and instance-to-instance (reference-based) rewards using our aesthetics model and CLAP score, respectively. As shown in Figure 3, when optimizing the aesthetics score, DRAGON consistently achieves at least 80% reward win rate over the baseline, validating that our aesthetics model encodes learnable information, and DRAGON has a strong optimization capability.
@@ 5.2 Optimizing Instance-Level Rewards – Predicted Aesthetics Score and CLAP Score
DPO versus KTO; paired versus unpaired. Using the aesthetics score as the reward function, we perform ablation studies on different loss functions described in Section 3.2. Specifically, we focus on DPO loss with paired demonstrations, KTO loss with paired demonstrations, and KTO loss with unpaired demonstrations. As shown in Figure 3, DPO-Paired slightly outperforms KTO-Paired. DPO sees more stable training, whereas KTO improves the model faster. While KTO is more flexible by allowing for unpaired demonstrations, KTO-Paired outperforms KTO-Unpaired. Hence, direct pair-wise contrasting signals are advantageous, and we should prefer paired demonstrations when available.
Ablation on diffusion steps. Reducing the number of diffusion steps significantly accelerates training-time online generation at the cost of demonstration quality, but how does this affect the model fine-tuned with DRAGON? As shown in Figure 3, reducing the training-time diffusion steps from 40 to 10 (but still generating test examples with 40 steps) only induces a tiny change in generation quality. If we also decrease the number of test-time inference steps to 10, then the model trained with 10-step demonstrations can even outperform the one trained with 40 steps. In all inference settings, both DRAGON models outperform the baseline, confirming the generalization of model improvement across diffusion step settings. Moreover, the DRAGON models have overall flatter quality-vs-steps curves, with their 10-step generations outperforming the baseline model’s 40-step generations. Hence, to achieve a similar generation quality, DRAGON can significantly reduce inference-time computation, manifesting some properties of distilled diffusion models such as consistency models (Song et al., 2023; Bai et al., 2024; Novack et al., 2025).
All subsequent experiments use the Diffusion-KTO loss function with paired 40-step demonstrations. Section B.3 presents additional analyses on aesthetics optimization.
@@ 5.3 Optimizing Instance-to-Distribution Reward – Per-Song FAD
We fine-tune our music generator to minimize per-song FAD. In Section C.1, we perform ablation studies to demonstrate the statistical correlation between per-song FAD and human aesthetic perception, providing a theoretical foundation for improving human-perceived music quality via optimizing per-song FAD. Table 1(a) presents the model performance when optimizing per-song FAD with different reference statistics.
@@
We first consider using audio embeddings of human-created music as the FAD reference statistics. As shown in Table 1(a), for CLAP and diffusion VAE encoders alike, when the reference is ALIM ground-truth embeddings, minimizing per-song FAD enhances the target reward as well as the aesthetics score. The VAE embeddings are particularly powerful – optimizing the per-song VAE-FAD (ALIM) not only achieves a 93.9% win rate in this metric, but also generalizes to multiple other metrics. Over all models, the improvement in the per-song FAD to ALIM statistics is highly correlated with the SDNV-FAD improvement. However, using SDNV statistics as the DRAGON optimization target is less effective, likely due to SDNV’s less consistent quality and smaller size, as well as the mismatch between SDNV FAD reference and ALIM training prompts. We thus highlight the importance of using high-quality reference music and avoiding dataset mismatches. Overall, these results show DRAGON can enhance music generation without human feedback.
Next, we leverage the cross-modality nature of the CLAP embedding space and use text embeddings as the FAD reference statistics for generated audio. Notably, by minimizing the per-song FAD to ALIM captions’ CLAP embeddings, DRAGON achieves all-around improvements across all metrics in Table 1(a). Surprisingly, the cross-metric generalization even outperforms optimizing FAD with audio reference. While it seems counter-intuitive that text can be more helpful than music, this result is explainable. As shown in Section C.1, compared to per-song FAD to audio reference, the audio-to-text FAD with ALIM captions is more strongly correlated with human preference. DRAGON can similarly optimize the per-song text-FAD to SDNV captions. Similar to the audio-reference case, when steering toward SDNV captions, while the reward win rate is high and the improvement generalizes to per-song audio-FAD, other metrics benefit less.
Optimizing per-song text-FAD is similar to optimizing CLAP score in that both achieve aesthetic improvements with ALIM captions only, without any ground-truth music. In nearly all metrics, optimizing per-song text-FAD outperforms optimizing CLAP score. Hence, our novel instance-to-distribution paradigm is more effective than the traditional instance-to-instance method. In conclusion, DRAGON can align music generation with human preference using high-quality captions without human-created music.
Given DRAGON’s capability of learning from text-only data, we experiment whether “ungrounded music descriptions” not associated with any music collection can also improve music generation. We specifically consider two sets of prompts: 800 human-created prompts in the DMA dataset, and 870 prompts generated by the Mixtral-8x7B large language model (LLM) (Jiang et al., 2024). To ensure LLM-generated prompt quality, we gather 10,506 initial prompts, and then use a greedy pruning algorithm similar to Algorithm 1 to minimize the dataset text-to-text CLAP-FAD with ALIM captions. As shown in Table 1(a), when using human-created prompts, which are noisy and not based on actual music, as the exemplar set, DRAGON extracts learnable information via per-song FAD optimization, achieving an 83.7% reward win rate while improving CLAP score and per-song CLAP-audio-FAD. Similarly, the LLM-created prompts also encode learnable information and generalize even better to other metrics. Hence, we conclude that DRAGON can learn from ungrounded text-only music descriptions.
@@ 5.4 Optimizing Distribution-to-Distribution Reward – Full-Dataset FAD
Next, we leverage DRAGON’s unique capability to learn from reward functions that evaluate distributions and minimize full-dataset FAD. We consider the same reference statistics as in the per-song FAD experiments in Section 5.3, and present the results in Table 1(b). Optimizing dataset FAD is a particularly challenging task due to reward ambiguity and strong starting point performance. Specifically, learning from a reward like $r_{\textrm{dist}}$ that evaluates a distribution is harder and noisier than learning from an instance-level $r_{\textrm{instance}}$ due to the ambiguity of each instance’s contribution to the reward. Meanwhile, because diffusion training implicitly matches the generated distribution to the true data, the baseline model already starts at a decent FAD.
@@
Despite these challenges, all but one dataset FAD optimization run improve the reward function they optimize as shown in Table 1(b). As with per-song FAD optimization, using the ALIM dataset as the reference statistics achieves multi-metric improvements, generalizing across FAD to different references and improving aesthetics score. This enhancement is observed for all three encoder-modality combinations (VAE, CLAP-audio, CLAP-text), with the CLAP text embeddings attaining the best performance. Moreover, via dataset FAD optimization, ungrounded text embeddings (Human-Text and Mixtral-Text) can enhance music generation, especially in terms of aesthetics score and CLAP score. On average, dataset FAD optimization achieves a reward win rate of 81.5%, matching the average improvement for per-song FAD optimization. We thus conclude that DRAGON can enhance diffusion models by directly minimizing dataset FAD.
@@
Compared to per-song FAD results, we observe several trends with dataset FAD optimization. First, VAE results are comparatively weaker and CLAP results are stronger. This makes sense because VAE embeddings are lower-dimensional, summarizing the entire generated distribution ${\mathcal{D}}_{\theta}$ with $\mu_{\theta}\in{\mathbb{R}}^{32}$ and $\Sigma_{\theta}\in{\mathbb{S}}_{+}^{32}$, totaling 1056 numbers. Hence, information is lost and optimization signals are weak. Second, cross-reference generalization is weaker. Improving per-song FAD to one reference statistic often means better per-song FAD to other references, a correlation not observed with dataset FAD. Recall that the references are always full-dataset even when the generated statistics are per-song. Hence, per-song FAD improvement may partially come from a per-song-versus-full-dataset gap shrinkage, which generalizes across reference statistics. This gap does not exist for dataset FAD, and hence imitating a distribution may imply moving away from others. Third, optimizing dataset FAD to SDNV often hurts other metrics. We believe this is because all training prompts are from ALIM, and hence using SDNV reference statistics induces confusion. Specifically, if we run DRAGON with ALIM prompts and ALIM reference statistics, then when tested on SDNV prompts, the FAD with respect to SDNV statistics improves (see Figure 10). However, if DRAGON pairs ALIM prompts with SDNV statistics, then we do not observe this improvement. In summary, with dataset FAD optimization, it is important to select a suitable encoder and a reference statistic that matches the prompt distribution.
@@
Additionally, optimizing dataset FAD preserves more generation diversity than optimizing instance-level rewards like aesthetics, CLAP score, and per-song FAD. Figure 4 shows that optimizing per-song FAD worsens Vendi diversity score by an average of 17.7%, whereas optimizing dataset FAD only loses 3.2%. The next section will show that DRAGON can also explicitly optimize Vendi, balancing aesthetics and diversity.
@@ 5.5 Optimizing Reference-Free Distributional Reward – Vendi Diversity Score
To demonstrate DRAGON’s capability to improve generation diversity, we select the CLAP embedding Vendi score as the reward function. During training, we compute the Vendi score across all demonstrations in each training batch. During evaluation, we compute Vendi score with all 2,185 test generations. As shown in Figure 4, when explicitly optimizing Vendi, DRAGON significantly increases the score, achieving a 40.84% relative improvement. However, since Vendi does not provide any music quality information, optimizing Vendi alone distorts the generations and hurts their aesthetics. To this end, we co-optimize Vendi score and aesthetics score by randomly selecting one of the two rewards at each training iteration with equal probability. The result is a model that simultaneously improves Vendi and aesthetics, producing diverse high-quality music. In summary, we find DRAGON can promote generation diversity.

### v2

@@ 5.1 Models, Datasets, Training Settings, and Evaluation Metrics
Baseline model and pre-training. We use the base diffusion model from Presto (Presto) to generate 32-second single-channel (mono) 44.1kHz audio. It includes a latent-space score-prediction de-noising module (LDM; EDM; songscore) based on DiT-XL (DiT) that takes in the noise level and a text embedding as conditioning signals, a convolutional variational autoencoder (VAE) that converts audio to and from the diffusion latent space (DAC), and a FLAN-T5-based text encoder (FLAN-T5). The baseline model is pre-trained with diffusion loss to convergence on a 3600-hour instrumental music dataset with musical-metadata-grounded synthetic captions, which we call the Adobe Licensed Instrumental Music dataset (ALIM). Our inference uses 40 diffusion steps with the second-order DPM sampler (DPM) with CFG++ ($w=0.8$) enabled in selected time steps (CFG++). Please see Appendix˜E for details.
Training and evaluation prompts. We use ALIM training prompts (same setting as in pre-training) for DRAGON fine-tuning. Our evaluation uses a combination of the captions in an independent ALIM test split (800 pieces), the captions in a non-vocal Song Describer subset (SongDescriber) (585 pieces, abbreviated as SDNV), and the real-world user prompts in the DMA dataset (800 pieces). Unless specified otherwise, all evaluation metrics are computed with generations from these 2,185 prompts (one generation per prompt).
@@
Correlation measures. In addition to evaluating and comparing DRAGON models, we offer correlation analyses between various reward signals. We quantify correlation with overall Pearson correlation (PLCC) and per-prompt Spearman’s rank correlation (SRCC) in Section˜B.2.
@@
FAD encoders. The per-song and dataset FAD use ALIM and SDNV as reference statistics. DRAGON training uses ALIM’s training split statistics to compute FAD, while evaluation uses the test split. We consider CLAP and the diffusion VAE as FAD encoders, with the former known to provide high-quality semantic embeddings (FADTK). To account for any VAE reconstruction inaccuracies, we decode the generated VAE embeddings to audio and back. Compared with CLAP, the VAE encoder produces considerably more embeddings per song (1,836 versus 9) in a much lower dimension (32 versus 512).
CLAP embedding setting. We use the music-specializing LAION-CLAP checkpoint (laionclap2023; HTS-AT). Since CLAP takes in 10-second 48kHz audio whereas our generations are 32-second 44.1kHz, pre-/post-processing is required. We focus on two settings: MA and FADTK. The MA setting is optimized for our music aesthetics predictor with ablation studies in Section˜B.2, while the FADTK setting follows (FADTK). More details about and comparisons between the two settings are in Section˜C.2.
Open-source comparison. We compare DRAGON with open-source models such as Stable Audio (StableAudio) and MusicGen (MusicGen) in Appendix˜D.
We first use DRAGON to optimize instance-wise (reference-free) and instance-to-instance (reference-based) rewards using our aesthetics model and CLAP score, respectively. As shown in Figure˜3, when optimizing the aesthetics score, DRAGON consistently achieves at least 80% reward win rate over the baseline, validating that our aesthetics model encodes learnable information, and DRAGON has a strong optimization capability.
DPO versus KTO; paired versus unpaired. Using the aesthetics score as the reward function, we perform ablation studies on different loss functions described in Section˜3.2. Specifically, we focus on DPO loss with paired demonstrations, KTO loss with paired demonstrations, and KTO loss with unpaired demonstrations. As shown in Figure˜3, DPO-Paired slightly outperforms KTO-Paired. DPO sees more stable training, whereas KTO improves the model faster. While KTO is more flexible by allowing for unpaired demonstrations, KTO-Paired outperforms KTO-Unpaired. Hence, direct pair-wise contrasting signals are advantageous, and we should prefer paired demonstrations when available.
Ablation on diffusion steps. Reducing the number of diffusion steps significantly accelerates training-time online generation at the cost of demonstration quality, but how does this affect the model fine-tuned with DRAGON? As shown in Figure˜3, reducing the training-time diffusion steps from 40 to 10 (but still generating test examples with 40 steps) only induces a tiny change in generation quality. If we also decrease the number of test-time inference steps to 10, then the model trained with 10-step demonstrations can even outperform the one trained with 40 steps. In all inference settings, both DRAGON models outperform the baseline, confirming the generalization of model improvement across diffusion step settings. Moreover, the DRAGON models have overall flatter quality-versus-steps curves, with their 10-step generations outperforming the baseline model’s 40-step generations. Hence, to achieve a similar generation quality, DRAGON can significantly reduce inference-time computation, manifesting some properties of distilled diffusion models such as consistency models (CM; ConsistencyTTA; Presto).
All subsequent experiments use the Diffusion-KTO loss function with paired 40-step demonstrations. Section˜B.3 presents additional analyses on aesthetics optimization.
@@ 5.3 Optimizing Instance-to-Distribution Reward – Per-Song FAD
We fine-tune our music generator to minimize per-song FAD. In Section˜C.1, we perform ablation studies to demonstrate the statistical correlation between per-song FAD and human aesthetic perception, providing a theoretical foundation for improving human-perceived music quality via optimizing per-song FAD. Table˜2(a) presents the model performance when optimizing per-song FAD with different reference statistics.
@@
We first consider using audio embeddings of human-created music as the FAD reference statistics. As shown in Table˜2(a), for CLAP and diffusion VAE encoders alike, when the reference is ALIM ground-truth embeddings, minimizing per-song FAD enhances the target reward as well as the aesthetics score. The VAE embeddings are particularly powerful – optimizing the per-song VAE-FAD (ALIM) not only achieves a 93.9% win rate in this metric, but also generalizes to multiple other metrics. Over all models, the improvement in the per-song FAD to ALIM statistics is highly correlated with the SDNV-FAD improvement. However, using SDNV statistics as the DRAGON optimization target is less effective, likely due to SDNV’s less consistent quality and smaller size, as well as the mismatch between SDNV FAD reference and ALIM training prompts. We thus highlight the importance of using high-quality reference music and avoiding dataset mismatches. Overall, these results show DRAGON can enhance music generation without human feedback.
@@
Next, we leverage the cross-modality nature of the CLAP embedding space and use text embeddings as the FAD reference statistics for generated audio. Notably, by minimizing the per-song FAD to ALIM captions’ CLAP embeddings, DRAGON achieves all-around improvements across all metrics in Table˜2(a). Surprisingly, the cross-metric generalization even outperforms optimizing FAD with audio reference. While it seems counter-intuitive that text can be more helpful than music, this result is explainable. As shown in Section˜C.1, compared to per-song FAD to audio reference, the audio-to-text FAD with ALIM captions is more strongly correlated with human preference. DRAGON can similarly optimize the per-song text-FAD to SDNV captions. Similar to the audio-reference case, when steering toward SDNV captions, while the reward win rate is high and the improvement generalizes to per-song audio-FAD, other metrics benefit less.
Optimizing per-song text-FAD is similar to optimizing CLAP score in that both achieve aesthetic improvements with ALIM captions only, without any ground-truth music. In nearly all metrics, optimizing per-song text-FAD outperforms optimizing CLAP score. Hence, our novel instance-to-distribution pa-radigm is more effective than traditional instance-to-instance. In conclusion, DRAGON can align music generation with human preference using high-quality captions without human-created music.
Given DRAGON’s capability of learning from text-only data, we experiment whether “ungrounded music descriptions” not associated with any music collection can also improve music generation. We specifically consider two sets of prompts: 800 human-created prompts in the DMA dataset, and 870 prompts generated by the Mixtral-8x7B large language model (LLM) (Mixtral). To ensure LLM-generated prompt quality, we gather 10,506 initial prompts, and then use a greedy pruning algorithm similar to Algorithm˜1 to minimize the dataset text-to-text CLAP-FAD with ALIM captions. As shown in Table˜2(a), when using human-created prompts, which are noisy and not based on actual music, as the exemplar set, DRAGON extracts learnable information via per-song FAD optimization, achieving an 83.7% reward win rate while improving CLAP score and per-song CLAP-audio-FAD. Similarly, the LLM-created prompts also encode learnable information and generalize even better to other metrics. Hence, we conclude that DRAGON can learn from ungrounded text-only music descriptions.
@@ 5.4 Optimizing Distribution-to-Distribution Reward – Full-Dataset FAD
Next, we leverage DRAGON’s unique capability to learn from reward functions that evaluate distributions and minimize full-dataset FAD. We consider the same reference statistics as in the per-song FAD experiments in Section˜5.3, and present the results in Table˜2(b). Optimizing dataset FAD is a particularly challenging task due to reward ambiguity and strong starting point performance. Specifically, learning from a reward like $r_{\textrm{dist}}$ that evaluates a distribution is harder and noisier than learning from an instance-level $r_{\textrm{instance}}$ due to the ambiguity of each instance’s contribution to the reward. Meanwhile, because diffusion training implicitly matches the generated distribution to the true data, the baseline model already starts at a decent FAD.
@@
Despite these challenges, all but one dataset FAD optimization run improve the reward function they optimize as shown in Table˜2(b). As with per-song FAD optimization, using the ALIM dataset as the reference statistics achieves multi-metric improvements, generalizing across FAD to different references and improving aesthetics score. This enhancement is observed for all three encoder-modality combinations (VAE, CLAP-audio, CLAP-text), with the CLAP text embeddings attaining the best performance. Moreover, via dataset FAD optimization, ungrounded text embeddings (Human-Text and Mixtral-Text) can enhance music generation, especially in terms of aesthetics score and CLAP score. On average, dataset FAD optimization achieves a reward win rate of 81.5%, matching the average improvement for per-song FAD optimization. We thus conclude that DRAGON can enhance diffusion models by directly minimizing dataset FAD.
@@
Compared to per-song FAD results, we observe several trends with dataset FAD optimization. First, VAE results are comparatively weaker and CLAP results are stronger. This makes sense because VAE embeddings are lower-dimensional, summarizing the entire generated distribution ${\mathcal{D}}_{\theta}$ with $\mu_{\theta}\in{\mathbb{R}}^{32}$ and $\Sigma_{\theta}\in{\mathbb{S}}_{+}^{32}$, totaling 1056 numbers. Hence, information is lost and optimization signals are weak. Second, cross-reference generalization is weaker. Improving per-song FAD to one reference statistic often means better per-song FAD to other references, a correlation not observed with dataset FAD. Recall that the references are always full-dataset even when the generated statistics are per-song. Hence, per-song FAD improvement may partially come from a per-song-versus-full-dataset gap shrinkage, which generalizes across reference statistics. This gap does not exist for dataset FAD, and hence imitating a distribution may imply moving away from others. Third, optimizing dataset FAD to SDNV often hurts other metrics. We believe this is because all training prompts are from ALIM, and hence using SDNV reference statistics induces confusion. Specifically, if we run DRAGON with ALIM prompts and ALIM reference statistics, then when tested on SDNV prompts, the FAD with respect to SDNV statistics improves (see Figure˜10). However, if DRAGON pairs ALIM prompts with SDNV statistics, then we do not observe this improvement. In summary, with dataset FAD optimization, it is important to select a suitable encoder and a reference statistic that matches the prompt distribution.
@@
Additionally, optimizing dataset FAD preserves more generation diversity than optimizing instance-level rewards like aesthetics, CLAP score, and per-song FAD. Figure˜4 shows that optimizing per-song FAD worsens Vendi diversity score by an average of 17.7%, whereas optimizing dataset FAD only loses 3.2%. The next section will show that DRAGON can also explicitly optimize Vendi, balancing aesthetics and diversity.
@@ 5.5 Optimizing Reference-Free Distributional Reward – Vendi Diversity Score
To demonstrate DRAGON’s capability to improve generation diversity, we select the CLAP embedding Vendi score as the reward function. During training, we compute the Vendi score across all demonstrations in each training batch. During evaluation, we compute Vendi score with all 2,185 test generations. As shown in Figure˜4, when explicitly optimizing Vendi, DRAGON significantly increases the score, achieving a 40.84% relative improvement. However, since Vendi does not provide any music quality information, optimizing Vendi alone distorts the generations and hurts their aesthetics. To this end, we co-optimize Vendi score and aesthetics score by randomly selecting one of the two rewards at each training iteration with equal probability. The result is a model that simultaneously improves Vendi and aesthetics, producing diverse high-quality music. In summary, we find DRAGON can promote generation diversity.